OpenStudy (anonymous):

A carpenter makes legs for 3-legged stools and 4-legged tables. He makes a total of 57 legs. He could have made legs for how many 3-legged stools? A. 8 B. 9 C. 10 D. 11

OpenStudy (cwrw238):

let there be x stools with 4 legs and y stools with 4 legs, then: 4x + 3y = 57 you now have to fit positive whole numbers to this equation

OpenStudy (cwrw238):

lets try 8 3 legged stools 4x + 3*8 = 57 4x = 57 - 24 = 33 this cant be correct as 33 is not a multiple of 4 do same with the other options

OpenStudy (cwrw238):

do you follow what i've written?

OpenStudy (anonymous):

yes thank you

OpenStudy (cwrw238):

try y = 11 ( 11 3-legged stools)